# Improving Local Training in Federated Learning via Temperature Scaling

### v2

@@ Appendix D Experiment on hyperparameters
In this section, we explore the effect of other federated learning hyperparameters (i.e., batch size, number of local epochs, number of participants, and degree of data disparity). Through this set of experiments, we observed that applying low temperatures during local training improves test accuracy and convergence speed, being independent from hyperparameters. We note that the experiments were done by applying the default setting except for the control factors.
D.1 Batch size
@@
Figure 10: Test accuracy and test loss over federated learning round with varying the batch size and the training temperature on the CIFAR10 dataset. Relatively small batch size shows better training quality with stability.
D.2 Number of participants
(a) Number of participants 5
(b) Number of participants 10
(c) Number of participants 15
Figure 11: Test accuracy and test loss over federated learning round with varying the number of participants and the training temperature on the CIFAR10 dataset. When relatively many clients participate in the training, it shows better training quality with stability
D.3 Number local epochs
@@
Figure 12: Test accuracy and test loss over federated learning round with varying the number of local epochs and the training temperature on the CIFAR10 dataset. Relatively many updates in the local training phase show better training quality with stability
D.4 Degree of data disparity
(a) $\alpha$=0.1
(b) $\alpha$=0.5
(c) $\alpha$=1.0
Figure 13: Test accuracy and test loss over federated learning round with varying the degree of data disparity and the training temperature on the CIFAR10 dataset. Relatively moderate data disparity shows better training quality with stability
Hyperparameters Appendix D total clients 50 communication rounds 100 optimizer SGD learning rate 0.001 local epoch 5 participants per round 5 batch size 32 $\alpha$ 0.5 loss function Cross Entropy
Table 7: Default hyperparameters used in this section. $\alpha$ denotes the importance of the Dirichlet distribution.
Appendix E Details of Centered Kernel Alignment (CKA)
E.1 Background
Centered kernel alignment (CKA) is a method to compute the similarity between two different neural networks by measuring the similarity of the network representation for the given input data. Let $X$ and $Y$ be the features generated from two different neural networks, we can calculate the CKA by Eq 3. As we are assuming a federated learning scenario, the clients share the same neural network architecture. Thus, the dimensions of the two features from each corresponding layer are identical, allowing for a direct comparison to be possible. We note that in our experiments, we exploited the linear CKA to measure the similarity of the feature spaces between two neural networks.
$\begin{split}CKA(X,Y)&=\frac{||Y^{T}X||^{2}_{F}}{||X^{T}X||_{F}||Y^{T}Y||_{F}}\\ \end{split}$ (3)
E.2 Limitation of Centered Kernel Alignment evaluation in federated learning
Here, we discuss our rationale for measuring CKA between a client model participating in the federated learning round and a reference model trained using i.i.d. configurations. As mentioned in Section 4.4, the feature space similarity itself only reports the similarity of neural networks and does not assure high quality (i.e., accuracy) of the trained models. Figure 14 illustrates the possible cases that can observed in a federated learning scenario. We use these four cases to understand how locally trained models converge to the reference model. Here, the orange line indicates the update directions of local models from the federated learning training phase (i.e., diverging from the initially shared aggregated model presented as the black dot) and the red dotted line represents the average of these locally updated models that occur at the server. The star represents the ideal model in the representation space with high accuracy.
(a) Case 1
(b) Case 2
(c) Case 3
(d) Case 4
(e) Case 5
Figure 14: Illustration of possible cases in a federated learning scenario. The solid orange line and dotted red line correspond to the local update direction and aggregated model update direction respectively. The black circle and the red star depict the initial point and optimal point respectively.
$\bullet$ High similarity between local models does not guarantee a better model (Case 1, 2): The first case arises when local models, each trained on their respective local data, produce highly similar updates due to a well-aligned data configuration. Such alignment implies that the training data across different clients are relatively similar in distribution and characteristics, allowing the aggregated global model to closely approach the ideal model achievable through centralized training. However, in non-i.i.d. environments, achieving this scenario is infrequent. In non-i.i.d. settings, the data distribution varies significantly across different local nodes. This variation can lead to local models learning disparate aspects of the data, resulting in updates that are less similar. Consequently, the aggregated global model may not perform as well as in an i.i.d. scenario, as it has to reconcile these diverse and potentially conflicting updates. On the other hand, Case 2 represents a critical scenario in federated learning where there is a high similarity among local model updates, but these updates are aligned in a suboptimal or incorrect direction. This situation can be considered one of the worst cases in federated learning. In addition, using these two cases, we demonstrate that the high similarity between local models does not ensure a superior update or quality of training.
$\bullet$ A superior model can be obtained without similar local models (Case 3): Case 3 illustrates a case where a superior model can be achieved even when local models exhibit relatively low similarity. In this scenario, each local model learns effective but diverse features from heterogeneous data configurations, and these diverse updates are properly aggregated at the central server. Unlike scenarios where local models produce highly similar updates, in this case, the updates from different local models show significant variation. This diversity reflects the different aspects of the data each local model has learned. This situation arises under conditions of data heterogeneity, where the data distributions across different clients are diverse. Each local model captures unique features relevant to its specific data subset. Despite the low similarity among local model updates, the central server effectively aggregates these diverse updates. The aggregation process leverages the complementary information from each local model, resulting in a robust and well-generalized global model. This case supports the hypothesis that a proper global model can be obtained even when local models do not exhibit high similarity and is a globally ideal case for federated learning scenarios.
$\bullet$ Relationship between model similarity and the magnitude of updates (Case 4, 5): It is important to note that the magnitude of updates can influence the perceived similarity of the network. If the update magnitude is small, the similarity of the network may be measured as high, whereas if the update magnitude is large, the similarity may be measured as relatively low. This example underscores the need to rethink the conventional emphasis on high similarity as an indicator of performance. It suggests that focusing solely on similarity metrics, while providing insights on how local training operations induce model changes, might overlook other critical factors that contribute to the effectiveness of the aggregated model. Therefore, a more nuanced approach to evaluating model updates and their impact on the accuracy performance is required, considering both the magnitude of updates and their alignment with the overall learning objectives.
Based on the aforementioned discussions, we introduce measuring similarity with a reference model as an alternative evaluation method in this work. Although we cannot ensure that the reference model trained in an i.i.d. data configuration perfectly represents the optimal update objective, we consider this method as an effective and reasonable way to evaluate the federated learning process. This approach allows for a more nuanced assessment of the updates by comparing them to a standard benchmark, helping to identify whether the aggregated model is progressing toward a semi-optimal solution despite the inherent data heterogeneity.
E.3 Experimental results
Table 8 presents the exact numbers for the accuracy and CKA values discussed in Section 4.4 and Figure 5. Figure 15 illustrates the data distribution applied in the experiments.
Temperature ($T$) Accuracy (%) Layer 1 CKA Layer 2 CKA Layer 3 CKA Layer 4 CKA 4.00 49.72 0.6539 0.5632 0.7149 0.7661 2.00 53.82 0.6792 0.6872 0.7417 0.7963 1.00 58.19 0.7419 0.6997 0.7644 0.8214 0.50 63.48 0.9322 0.7459 0.8181 0.8559 0.25 66.85 0.9569 0.8161 0.8418 0.8830 0.05 72.76 0.9390 0.9059 0.8898 0.9026 Reference Model ($T$=1.00, i.i.d.) 73.86 1.0000 1.0000 1.0000 1.0000
Table 8: Accuracy and CKA values of 2 layer CNN measured on the CIFAR10 dataset in Figure 5.
Figure 15: Visualization of the distribution of training data used in CKA evaluation. Best viewed in color.
(a) $T$=0.05
(b) $T$=0.25
(c) $T$=0.5
(d) $T$=1.0
(e) $T$=2.0
(f) $T$=4.0
Figure 16: Client-wise feature space similarity matrix measured in the first convolution layers (Layer 1) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100). (INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
(a) $T$=0.05
(b) $T$=0.25
(c) $T$=0.5
(d) $T$=1.0
(e) $T$=2.0
(f) $T$=4.0
Figure 17: Client-wise feature space similarity matrix measured in the second convolution layers (Layer 2) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100). (INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
(a) $T$=0.05
(b) $T$=0.25
(c) $T$=0.5
(d) $T$=1.0
(e) $T$=2.0
(f) $T$=4.0
Figure 18: Client-wise feature space similarity matrix measured in penultimate layers (Layer 3) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100). (INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
(a) $T$=0.05
(b) $T$=0.25
(c) $T$=0.5
(d) $T$=1.0
(e) $T$=2.0
(f) $T$=4.0
Figure 19: Client-wise feature space similarity matrix measured in classification layers (Layer 4) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100). (INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)

### v3

@@ 2. Global aggregation:
In this section, we explore the effect of other federated learning hyperparameters (i.e., number of participants (Figure 18), batch size (Figure 19), and number of local epochs (Figure 20)). Through this set of experiments, we observed that applying low temperatures during local training improves test accuracy and convergence speed, being independent of hyperparameters. We note that the experiments were done by applying the default setting except for the control factors.
One can view participant count, batch size, and local epochs through the lens of gradient variance. Sampling more clients per round increases the chance that the aggregated update better approximates the true gradient, reducing variance. Likewise, using smaller batch sizes with more local epochs yields more frequent updates that capture finer-grained information from each client’s data, which also dampens variance. Because low temperatures can sometimes introduce training instability, tuning these variance-reducing parameters, more clients, smaller batches, or additional local epochs may help stabilize convergence at low temperatures.
(a) Number of participants 5
(b) Number of participants 10
(c) Number of participants 15
Figure 18: Test accuracy and test loss over federated learning rounds with varying the number of participants and the training temperature on the CIFAR10 dataset. The shaded areas indicate the standard deviation across different runs. When relatively many clients participate in the training, it shows better training quality with stability.
@@
Figure 19: Test accuracy and test loss over federated learning rounds with varying the batch size and the training temperature on the CIFAR10 dataset. The shaded areas indicate the standard deviation across different runs. A relatively small batch size shows better training quality with stability.
@@
Figure 20: Test accuracy and test loss over federated learning rounds with varying the number of local epochs and the training temperature on the CIFAR10 dataset. The shaded areas indicate the standard deviation across different runs. Relatively many updates in the local training phase show better training quality with stability.
$\bullet$ Extended Stability Analysis Under Long-Horizon Training
To further assess the stability of low-temperature training beyond the main experimental horizon, we conducted an extended study with 1,000 federated training rounds. This experiment aims to evaluate whether the aggressive updates induced by Logit Chilling remain stable under prolonged optimization.
Figure 21 presents the convergence trajectories for multiple temperature settings. Across all configurations, models trained with lower temperatures continue to exhibit stable convergence behavior. Notably, temperatures below one consistently outperform the standard/high temperature setting ($T\in\{1.0,2.0,4.0\}$) throughout the extended training horizon, reaffirming the robustness and long-term benefits of applying temperature scaling during local training. These results demonstrate that the accelerated convergence observed in the main experiments does not introduce long-term instability, even when the training duration is substantially increased.
Figure 21: Average global test accuracy and test loss for 1,000 federated training rounds across different temperature settings. Lower temperatures maintain stable convergence and continue to outperform the standard/high temperature ($T\in\{1.0,2.0,4.0\}$).
$\bullet$ Analysis on Model Calibration
Figure 22: Reliability diagrams illustrating the calibration performance of models trained with different temperatures. Each diagram compares predicted confidence against empirical accuracy, showing that temperature scaling during training does not introduce notable miscalibration.
Figure 23: Prediction confidence distributions of models trained with varying temperatures. The histograms show how the confidence levels are distributed across samples, providing insight into potential over- or under-confidence behaviors under different temperature settings.
Figure 24: Expected Calibration Error (ECE) for models trained with different temperatures. Error bars indicate standard deviation across runs. The results demonstrate minimal differences in calibration quality across temperature settings.
Since temperature scaling was originally introduced for model calibration, one might be concerned that training with low temperatures could lead to over- or under-confidence. Model calibration measures how well a model’s predicted probabilities correspond to the true likelihood of correctness. A well-calibrated model outputs high confidence only when its predictions are likely to be correct, and expresses appropriate uncertainty when the outcome is less certain. Thus, calibration evaluates not only what the model predicts but also how confident it is in those predictions. For example, a model that frequently produces high-confidence predictions but achieves only moderate accuracy is considered over-confident, whereas a model that predicts with low confidence despite high accuracy is under-confident.
A standard quantitative metric for evaluating calibration is the Expected Calibration Error (ECE), defined as
$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right|,$
where $B_{m}$ denotes the set of samples whose predicted confidences fall into the $m$-th bin, $\mathrm{acc}(B_{m})$ is the empirical accuracy for that bin, $\mathrm{conf}(B_{m})$ is the average predicted confidence, and $n$ is the total number of samples. Lower ECE values indicate better calibration.
To assess calibration quality in our setting, we examine the prediction probability distributions, the model’s average confidence, and its accuracy. Concretely, we present the calibration results (Figure 22), the prediction confidence distributions (Figure 23), and the corresponding ECE values (Figure 24) for models trained with different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0). We note that we use the same temperature for testing as for training. Figure 23 shows that models trained with higher temperatures produce fewer samples in the high-confidence region, which is consistent with their lower performance at the same training round and indicates that they are less certain about the given inputs. In contrast, models trained with lower temperatures assign higher confidence to a larger portion of samples, and Figure 22 shows that these high-confidence regions remain well calibrated: the empirical accuracy in the highest confidence bins does not deteriorate, and the reliability curves stay close to the diagonal. Finally, as shown in Figure 24, across our experimental settings, we do not observe meaningful differences in ECE among models trained with different temperatures. Taken together, the ECE values, confidence distributions, and reliability diagrams indicate that applying temperature scaling during training does not introduce notable calibration degradation, nor systematic over-confidence or under-confidence, in our setting.
$\bullet$ Comparison to Knowledge Distillation
In this section, we clarify how our use of temperature scaling in FLex&\&Chill differs fundamentally from conventional knowledge distillation techniques. Although both rely on a temperature parameter applied to the softmax function, their goals, where the temperature is applied, and the corresponding training regimes are distinct.
In classical knowledge distillation, a high-capacity teacher model is first trained in a centralized manner. A temperature value larger than 1 is then applied to the teacher’s logits to produce softened probability distributions. These softened probabilities are used as soft labels to train a student model, typically by minimizing a distillation loss such as the Kullback–Leibler divergence between the teacher and student outputs, often combined with a standard cross-entropy loss on hard labels. The key role of temperature in this setting is to smooth the teacher’s output distribution so that relative relationships among non-argmax classes become more informative to the student. Importantly, in conventional knowledge distillation, the temperature is introduced primarily to generate softened targets and to define the distillation loss: both the teacher and the student outputs are typically divided by the same temperature $T>1$ inside the softmax when computing the distillation term, while the hard-label loss and inference-time predictions of the student still use temperature $T=1$. In other words, the temperature is mostly confined to the soft-label matching objective, rather than being used as a global training knob that systematically controls all aspects of the student’s optimization dynamics.
In contrast, our work does not involve a teacher–student architecture or soft-label transfer. FLex&\&Chill applies temperature scaling directly to the softmax of the model that is being trained on each client, and this is done throughout the federated training process rather than as a post-training calibration or label-smoothing step. Moreover, we operate in the low-temperature regime where the temperature lies between 0 and 1, which sharpens the output distribution and, as analyzed in our convergence theory, amplifies gradient magnitudes and accelerates optimization in the presence of non-i.i.d. data. In other words, temperature in our setting is used as an optimization control knob that shapes gradient flow and feature-space alignment across heterogeneous clients, not as a mechanism for compressing or transferring knowledge from a pre-trained teacher network. These differences imply that FLex&\&Chill is complementary to, rather than a variant of, conventional knowledge distillation.